Hole 1107A
The reentry cone was positioned in the center of the moonpool doors. The 48.8-m 16-in casing string was made up and latched into the reentry cone in ~5 hr. Making up the 16-in casing was slowed by excessive roll caused by two large cross swells affecting vessel stability. The jetting BHA was made up of a 14-3/4-in bit, bit sub, five 8-1/4-in drill collars, 16-in Dril-Quip CADA casing running tool, two 8-1/4-in drill collars, tapered drill collar and two stands of 5-1/2-in drill pipe. The jetting BHA was latched into the 16-in casing hanger and the reentry cone/16-in casing string assembly was lowered to the seafloor. The VIT was also deployed to observe the position of the reentry cone after jetting in, and to observe release of the 16-in casing running tool.

No pilot hole or jet test was performed prior to jetting in the 16-in casing at Hole 1107A. The location of Hole 1107A was positioned between Holes 757B and 757C which are 200 m apart and were drilled on Leg 121. The length of 16-in casing to be jetted-in at Hole 1107A was determined based on advanced hydraulic piston corer data from Leg 121. Also, the seafloor depth established during Leg 121 was used for starting the jetting in process for Hole 1107A. However, the official water depth of 1659 mbrf was determined by tagging the seafloor with the end of the 16-in casing, which was determined to occur at the first noticeable pressure change while initiating the jetting in process.

Jetting in of the 16-in casing began at 1455 hr 23 May, establishing Hole 1107A. Jetting in the 48.8-m of 16-in casing was accomplished in 2 hr and 21 min. The VIT was used to determine that the reentry cone mud mat was positioned properly, sitting on the seafloor. The 16-in casing running tool was unlatched and retrieved with no difficulties. During the trip out of the hole, two Woods Hole Oceanographic Institution (WHOI) ocean-bottom seismometers (OBSs) were deployed and surveyed relative to Hole 1107A as part of a seismic-while-drilling (SWD) experiment. A ship-to-ship transfer from the Sonne of blasting caps and a ranging transducer was also completed.

Once back on deck, the 16-in casing running tool was removed and an additional 8-1/4-in drill collar was added to make up the 14-3/4-in drilling BHA. The 14-3/4-in drilling BHA was tripped to the seafloor and the reentry cone reentered at 0543 hr 24 May. The hole was drilled ahead to a depth of 2081 mbrf (422 mbsf) with mud sweeps as follows; 1 x 25 bbl at 1810.4 mbrf (151.4 mbsf), 1 x 25 bbl at 1897 mbrf (238.1 mbsf), 1 x 30 bbl at 2081 mbrf (422 mbsf). As soon as the BHA was below the seafloor the Lamont-Doherty Earth Observatory (LDEO) sensor sub, part of the SWD experiment, was installed. The sediment portion of the hole drilled slower than anticipated with rates of penetration (ROPs) as low as 4 m/hr. However, the hole remained stable with little if any fill at connections.

Contact with basement was made at a depth of ~2030 mbrf (371 mbsf). Unlike the sediments, the basement initially drilled faster than anticipated with ROPs as high as 8 m/hr. The original drilling plan called for ending the 14-3/4-in hole at a depth of 2069 mbsf (410 mbsf). However, firm basement was not encountered until a depth of 2063 mbrf (404 mbsf) had been reached. Therefore, the decision was made to continue drilling ahead until at least one full joint of casing could be positioned below the hard basement contact.

A wiper trip was made to a depth of 1692 mbrf (33 mbsf, inside the 16-in casing) with a maximum overpull of 10,000 lb. The LDEO sensor sub was removed during the wiper trip in preparation of running the 10-3/4-in casing string. The hole was allowed to stabilize for one hour, and then the bit was slowly tripped back to bottom. A moderate drag of 20,000 lb was observed near the sediment/basement contact at a depth of 2035 mbrf (376 mbsf). The top drive was picked up, and the hole was washed and reamed through the sediment/basement contact area. Once the bit was back on bottom, the hole was swept with a 30-bbl mud pill and then displaced with 346 bbl of sepiolite mud. The bit was then retrieved in preparation for deploying the 10-3/4-in casing string.

A 413.2-m-long string of 10-3/4-in casing, including the remaining 23 joints of 10-3/4-in standard coupling casing on board, was made up in ~5 hr. A cementing BHA consisting of a 10-3/4-in SSR cementing plug, a 10-3/4-in Dril-Quip casing running tool, five 8-1/4-in drill collars, a tapered drill collar, and two stands of 5-1/2-in drill pipe was made up and latched into the 10-3/4-in in casing string. The 10-3/4-in casing string was then tripped to the seafloor, and Hole 1107A was reentered at 1515 hr 26 May.

The casing was run to depth when the top drive, kelly cock valve (used to prevent cement from flowing up into the top drive), and cementing head were picked up. The drill string was then spaced out to land the 10-3/4-in casing hanger inside the 16-in casing hanger. After landing the 10-3/4-in casing hanger, latch-in was confirmed with a 10,000-lb overpull. From reentry to latch-in confirmation took ~2 hr, with no difficulties encountered.

The 10-3/4-in casing was successfully cemented in place with 238 sacks (48.8 bbl) of blended 15.8 lb/gal cement. While displacing the cement, the SSR plug dart was dropped on top of the cement slug and chased with a 10-bbl fresh water spacer and then seawater. Landing of the SSR plug dart was confirmed with an increase in drill string pressure. While we pressured up the drill string to release the SSR plug, the cementing hose burst at ~2000 psi, requiring the pump to be shut down immediately. Fortunately, at that time the cement slug was contained within the casing below the SSR plug (with latched in dart) and the casing shoe.

Using a tugger, the driller was raised in the derrick to the heaving cementing head, where he broke off the cementing hose and replaced it with a blanking plug. Once this had been accomplished, the driller also had to manually open the kelly cock valve above the cementing head, so a circulation path through the top drive could be established. After the driller was safely back on the rig floor, the pump was re-engaged and the drill string pressured up immediately, indicating the SSR plug and dart had remained in place. The drill string pressure was increased to ~2800 psi when the SSR plug released and was pumped to the casing shoe. Upon landing of the SSR plug in the casing shoe, a solid 600-psi pressure was held for a few minutes. The stand pipe bleed-off valve was then opened and no flow back was observed, indicating the casing shoe valve was holding. The only task remaining was to release the casing running tool. It required ~45 min of working the 10-3/4-in casing running tool before it finally was released and the drill string could be retrieved.

The original drilling plan called for coring Hole 1107A to a depth of ~200-m into basement. However, with the lost time in port rebuilding the lower guide horn, there was not time to core at all, let alone hope of reaching a depth of 200 m into basement. The decision was made to deploy a 9-7/8-in tricone drill bit and drill as deep into basement as the remaining operational time allowed in the hope of leaving an acceptable hole for installation of an ION seismometer in the future. A 9-7/8-in drilling BHA was made up consisting of a 9-7/8-in tricone bit, bit sub, eight 8-1/4-in drill collars, a tapered drill collar, and two stands of 5-1/2-in drill pipe. The 9-7/8-in drilling BHA was tripped to the seafloor, and Hole 1107A was reentered at 0738 hr 27 May.

The 10-3/4-in casing shoe was drilled out in 43 min. The rathole below the 10-3/4-in casing shoe was cleaned out, and then a 9-7/8-in diameter hole was drilled into basement to a total depth (TD) of 2152.8 mbrf (493.8 mbsf, 122.8 m into basement, 79.8 m below the 10-3/4-in casing shoe). A schematic drawing of the NERO borehole installation is shown in Figure 27. A single 25-bbl mud sweep was pumped at 2107.4 mbrf (448.4 mbsf). At TD the hole was flushed clean and a 30-bbl mud sweep was circulated. With the hole cleaned up as much as possible, and with operations time expiring, the bit was retrieved. The bit cleared the rotary table at 0640 hr 28 May, ending Hole 1107A. During the trip out of the hole, the two WHOI OBSs were recovered. The vessel was secured for sea and the transit to Darwin, Australia, began at 0730 hr 28 May.

To 179 OPS Figure Captions

To 179 Table of Contents